OpenStudy (anonymous):

7w-3 ------ +3<5 2 how can I find the solution set?

OpenStudy (anonymous):

\[\left|7w-3 \right|\]

OpenStudy (anonymous):

The 7w-3 is an absoulte value

OpenStudy (anonymous):

Well, start by moving the 3, to the other side and clearing the fractional part.

OpenStudy (anonymous):

3+7w-3=5 right?

OpenStudy (anonymous):

3+7w-3<5 I mean

OpenStudy (anonymous):

No. \[\frac{|7w-3|}{2} + 3 < 5\] \[\implies \frac{|7w-3|}{2} < 5 -3\] \[\implies |7w-3|< 2(5 + 3)\] Right?

OpenStudy (anonymous):

Ack!

OpenStudy (anonymous):

Should be \[|7w -3| < 2(5-3)\]

OpenStudy (anonymous):

7w-3<22

OpenStudy (anonymous):

Did you get 7w<22/3?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

2*2 = 4.

OpenStudy (anonymous):

That means that 7w is less than 4 units away from 3.

OpenStudy (anonymous):

The part of the equation where you have 7w-3<2(5+3) =7w-3<10+6

OpenStudy (anonymous):

I corrected that. It should be 2(5-3)

OpenStudy (anonymous):

Not 2(5+3)

OpenStudy (anonymous):

So that means -4 < 7w - 3 < 4

OpenStudy (anonymous):

Your right. And from there 7w<4-3 =7w<1 =w<1/7?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

\[|7w - 3| < 4 \]\[\implies -4 < 7w -3 < 4\]

OpenStudy (anonymous):

Then you add 3 to each part of the relation.

OpenStudy (anonymous):

\[-4 + 3 < 7w - 3 + 3 < 4 + 3\]

OpenStudy (anonymous):

Wow thats alot

OpenStudy (anonymous):

Think about it. If I tell you that | x | < 1, That means that x is somewhere between 1 and -1.

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

So if I tell you that |x + 3| < 4, that means that x+3 is somewhere between -4 and 4

OpenStudy (anonymous):

So now we simplify and we have: \[-1 < 7w < 7\]

OpenStudy (anonymous):

ok, so I thought when we were at the 7w-3<4, we were almost finished. But we have to bring the 4 to the other side also and do it again.

OpenStudy (anonymous):

well you can work the two relations at the same time.

OpenStudy (anonymous):

Dividing the whole thing by 7 we get...?

OpenStudy (anonymous):

1/7<w

OpenStudy (anonymous):

-1/7 is part of it. what's the other part?

OpenStudy (anonymous):

-1/7<w<7/7

OpenStudy (anonymous):

Right.

OpenStudy (anonymous):

7/7=1

OpenStudy (anonymous):

-1/7<w<1 So would it be accurate that the solution set would become (-1/7, 1)?

OpenStudy (anonymous):

So w can be anything between -1/7 and 1 and the original relation \[\frac{|7w -3|}{2}+3 < 5\] Will be true.

OpenStudy (anonymous):

Yep. (-1/7,1) is the solution set.

OpenStudy (anonymous):

ok, great. I understand now. Thanks for your help.

OpenStudy (anonymous):

Just remember when you take off the absolute values that you have 2 possible solutions. \[|x| = a \implies x \in \{a,-a\}\] \[|x| < a \implies -a < x < a \] \[|x| \le a \implies -a \le x \le a \] \[|x| \ge a \implies x \le -a \ or\ a\le x \]

OpenStudy (anonymous):

etc.

OpenStudy (anonymous):

Would the technique be the same if the absolute value contained a fraction? -2\[-2\left| 7-v/8 \right|+3-11?\] First step move the 8 over?

OpenStudy (anonymous):

\[2\left| 7-v/8 \right|\]

OpenStudy (anonymous):

\[-2\left| 7-v/8 \right|-3\]

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

Sorry, was cleaning

OpenStudy (anonymous):

Yes, it'd be the same technique move everything that's not inside the absolute value to the other side, then evaluate the absolute value by splitting it up into parts.